white

1

white

2

red

3

red

4

red

5

white

6

black

7

red

8

red

9

red

10

black

11

Here starts hyperlink: CTAN ended. TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG TUG Again starts hyperlink: CTAN all done!